A 108.2 mL sample of 1.00 M NaOH is mixed with 54.1 mL of 1.00 M H2SO4 in a large Styrofoam coffee cup; the cup is fitted with a lid through which passes a calibrated thermometer. The temperature of each solution before mixing is 23.05 °C. After adding the NaOH solution to the coffee cup and stirring the mixed solutions with the thermometer, the maximum temperature measured is 30.50 °C. Assume that the density of the mixed solutions is 1.00 g/mL, that the specific heat of the mixed solutions is 4.18 J/(g·°C), and that no heat is lost to the surroundings.

Calculate the enthalpy change per mole of H2SO4 in the reaction.
Answer #1

Given, m= 108.2 + 54.1 = 162.3 moog i density 1.00g/ml) Cp- 4.184 319 0 AT= 30.50 - 23.05 = 7.45 . At = mc AT = 162.344.184 xthank you vmuch.....